OpenStudy (samslays):

HELP!! WILL FAN + MEDAL Determine whether the equation represents a direct variation. If it does, find the constant of variation 2y=5x+1

OpenStudy (samslays):

@Susanna

OpenStudy (susanna):

y=kx is a direct variation.

OpenStudy (susanna):

2y=5x+1 y=(5/2)x+1/2

OpenStudy (samslays):

I know the equation used to find the constant, but I am a little confused on whether it can represnt a direct variation equation with the " +!" at the end

OpenStudy (susanna):

I think it represents a direct variation. The constant of the variation is 5/2

OpenStudy (samslays):

@Susanna could you explain to me how you got 5/2? I am getting really confused

OpenStudy (pawanyadav):

You have to get y on LHS and remaining in RHS So 2y=5x+1 Divide by 2 both side to get y only in LHS So y=(5/2)x+(1/2)

OpenStudy (pawanyadav):

@Samslays